OpenStudy (anonymous):

solve the equation for y: 4x+3y=-12

OpenStudy (anonymous):

y=-(12+4x)/3

OpenStudy (anonymous):

4x=-12-3y x=-3-3/4y

OpenStudy (anonymous):

matticked please explain how you got the answer